Too long since my last post.
* GMOS pre-imaging re-oriented, re-submitted, re-activated. We'll lose 10 minutes from the science exposures, but will still have 2 minutes' more than we asked for! Still need to take a stab at mask design.
* Started a push to cross-correlate all available, good SDSS quasar spectra with the photometric etc. info of their targets, for future reference and so Ali Rafiee can reconstruct the spectra. Progress made, but couldn't finish by self-imposed deadline, so set it aside for higher-priority project. Will return to it tonight or tomorrow.
* Risaliti et al. paper on [O III] EWs for quasar club today; a worthwhile paper.
* Monday and Tuesday and part of Wednesday was selection of more PHL 1811 analogs for next Chandra AO. Target list seems to have come together nicely, although I am now wondering again about O VI BAL quasars... which may be good for a proposal!